Prove that Z[√-3 ] is not a U.F.D
1
Expert's answer
2020-12-01T15:04:48-0500

Let consider the number 4 from "Z[\sqrt{-3} ]"


"4 = 2*2 = (1+\sqrt{-3}) *(1 - \sqrt{-3})"


So 4 is not uniquely factorized in "Z[\sqrt{-3} ]" .


Hence "Z[\sqrt{-3} ]" is not a U.F.D (Unique Factorization Domain)